40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 29 20:21:29 2016, Johannes, Update, General, YARM loss measurement 
    Reply  Mon Oct 3 21:24:02 2016, Johannes, Update, General, XARM loss measurement ReflectionLoss.pdf
       Reply  Tue Oct 4 22:18:24 2016, Johannes, Update, General, X/YARM loss measurement anomalousData.png
          Reply  Wed Oct 5 19:10:04 2016, gautam, Update, General, Arm loss measurement review 
             Reply  Mon Nov 14 19:15:57 2016, Johannes, Update, General, Achievable armloss measurement accuracy 
                Reply  Tue Nov 15 20:35:19 2016, Johannes, Update, General, Achievable armloss measurement accuracy 
                   Reply  Thu Nov 17 21:54:11 2016, Johannes, Update, General, Achievable armloss measurement accuracy 
                      Reply  Thu Jan 12 02:45:53 2017, Johannes, Update, General, Next armloss steps ass_illustration.pdf
                         Reply  Fri Jan 13 08:54:32 2017, Johannes, Update, General, DC PD installed ASDCPD_up.jpgASDCPD_down.jpgscrambled_osci.jpg
Message ID: 12710     Entry time: Fri Jan 13 08:54:32 2017     In reply to: 12704
Author: Johannes 
Type: Update 
Category: General 
Subject: DC PD installed 

I installed a DC PD (Thorlabs PDA 520) in the beam path to AS55. I placed a 2" 90/10 BS on a flip mount that picks of enough light for the PD to spit out ~8V when the port is bright. Single arm continuous signal will be ~2V. While most of the light still continues towards AS55, the displacement from the BS moves the beam off AS55, so I used the flip mount in case anyone needs to use AS55. The current configuration is UP.

When we're done with loss investigations the flip mount should be removed from the bench.

I hooked the PD up to an ethernet-enabled scope and started scripting the loss map measurement (scope can receive commands via http so we can automate the data acquisition). The scope that was present at the bench and had been used for the MC ringdown measurements had a 'scrambled' screen that I couldn't fix so I had to retrieve another scope ("scope1"). I'll try to find out what's wrong with it but we may have to send it in for repair.

 

ELOG V3.1.3-